Solve the linear systems usung any method:
Eq.1 : 1/3x - 3/4y = -1/2
Eq.2 : 1/6x + 1/8y = 3/4

Answer by edjones(8007)

Eq.1 : 1/3x - 3/4y = -1/2
Eq.2 : 1/6x + 1/8y = 3/4
.
Get rid of the fractions!
-4x+9y=6 multiply times -12
4x+3y=18 multiply times 24
-----------add
==>12y=24
y=2
4x+6=18 substitute 2 for y in Eq. 2
4x=12
x=3
.
Ed